# Nyström Kernel Stein Discrepancy

### v1

@@ Appendix D External statements
This section collects the external statements that we use. Lemma D.1 gives equivalent norms for $f\otimes f$. We collect properties of Orlicz norms in Lemma D.2. Theorem D.3 is about the concentration of the empirical covariance, and Theorem D.4 recalls the Bernstein’s inequality for separable Hilbert spaces. Theorem D.5 is a concentration result for bounded random variables in a separable Hilbert space. Theorem D.6 recalls the combination of Rudi et al. [2015, Proposition 3 and Proposition 7].
@@
Define the sampling operator $Z_{m}:\mathcal{H}_{h_{p}}\to\mathbb{R}^{m}$ by $f\mapsto\frac{1}{\sqrt{m}}\left(f(\tilde{\mathbf{x}}_{i})\right)_{i=1}^{m}$. Its adjoint $Z^{*}_{m}:\mathbb{R}^{m}\to\mathcal{H}_{h_{p}}$ (see Sterge and Sriperumbudur [2022, Lemma A.7(i)] is given by $\bm{\alpha}=\left(\alpha_{i}\right)_{i=1}^{m}\mapsto\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\alpha_{i}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)$. Recall that $\mathcal{H}_{h_{p},m}=\mathrm{span}\left\{h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\mid i\in[m]\right\}$ and notice that $\operatorname{range}P_{\mathcal{H}_{h_{p}}}=\overline{\operatorname{range}Z^{*}_{m}}$. We obtain
$\left\|{\left(I-P_{h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\lambda\left\|{\left(Z^{*}_{m}Z_{m}+\lambda I\right)^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(b)}}{{=}}\lambda\left\|{\tilde{C}_{\hat{Q}_{n},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}$ (D.69)
@@
$\lambda\left\|{\tilde{C}_{\hat{Q}_{n},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)},$ (D.70)

### v2

@@ Appendix D External statements
This section collects the external statements that we use. Lemma D.1 gives equivalent norms for $f\otimes f$. We collect properties of Orlicz norms in Lemma D.2. Theorem D.3 is about the concentration of the empirical covariance, and Theorem D.4 recalls Bernstein’s inequality for separable Hilbert spaces. Theorem D.5 is a concentration result for bounded random variables in a separable Hilbert space. Theorem D.6 recalls the combination of Rudi et al. [2015, Proposition 3 and Proposition 7].
@@
Define the sampling operator $Z_{m}:\mathcal{H}_{h_{p}}\to\mathbb{R}^{m}$ by $f\mapsto\frac{1}{\sqrt{m}}\left(f(\tilde{\mathbf{x}}_{i})\right)_{i=1}^{m}$. Its adjoint $Z^{*}_{m}:\mathbb{R}^{m}\to\mathcal{H}_{h_{p}}$ (see Sterge and Sriperumbudur [2022, Lemma A.7(i)] is given by $\bm{\alpha}=\left(\alpha_{i}\right)_{i=1}^{m}\mapsto\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\alpha_{i}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)$. Recall that $\mathcal{H}_{h_{p},m}=\mathrm{span}\left\{h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\mid i\in[m]\right\}$ and notice that $\operatorname{range}P_{\mathcal{H}_{h_{p},m}}=\overline{\operatorname{range}Z^{*}_{m}}$. We obtain
$\left\|{\left(I-P_{h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\lambda\left\|{\left(Z^{*}_{m}Z_{m}+\lambda I\right)^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(b)}}{{=}}\lambda\left\|{\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}$ (D.60)
@@
$\lambda\left\|{\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)},$ (D.61)